Use Substitutionto solve each euqations, If the system does not have exactly one solution, state whethter it has no solution or infinitely many solutions.
1/3X-Y=3 2X+Y=25

Answer by Alan3354(69443)

Use Substitutionto solve each euqations, If the system does not have exactly one solution, state whethter it has no solution or infinitely many solutions.
1/3X-Y=3 2X+Y=25
-------------------
1/3X-Y=3 --> x-3y = 9 --> x = 3y+9
2X+Y=25
Sub for x into 2nd eqn
2(3y+9) + y = 25
6y+18 + y = 25
7y = 7
y = 1
-------------
Sub for y into either eqn
2x + 1 = 25
2x = 24
x = 12